OpenStudy (anonymous):

A rock has 12.5% of its original amount of potassium-40 remaining in it; potassium-40 has a half-life of 1.25 billion years. How long ago was the rock formed?

OpenStudy (aaronq):

do you know how to do these?

OpenStudy (anonymous):

I am horrible at these problems. I would love some help!

OpenStudy (aaronq):

i'll show you the proper way, which you can solve any odd amount of time or amount of substance. These are first order reactions so you can use this to find the decay constant, k: \[t _{1/2}=\frac{ \ln2 }{ k }\] t1/2= half life then plug it into the exponential decay/growth formula: \[A=A _{0}e ^{-kt}\] Ao= initial amount A= amount after time elapsed t = time k = decay constant

OpenStudy (anonymous):

What is the decay constant here?

OpenStudy (aaronq):

you find it through the first equation

OpenStudy (anonymous):

-.94

OpenStudy (aaronq):

how are you doing that?

OpenStudy (anonymous):

On my Ti-83, that's what I got for k.

OpenStudy (aaronq):

no i mean, what did you do?

OpenStudy (anonymous):

I divided 1.25 by 2, and then multipled it by LN, and multiplied that by 2.

OpenStudy (aaronq):

thats not how you solve it, 1.25 = ln2/k right? isolate k, k=ln2/1.25

OpenStudy (aaronq):

ln2 is a number, don't take it separately, it's approx 0.693

OpenStudy (anonymous):

0.47?

OpenStudy (anonymous):

Wow, I was really off. Thank you for setting me straight.

OpenStudy (aaronq):

it's 0.5545

OpenStudy (anonymous):

I didn't take it separately..

OpenStudy (anonymous):

So, how do I find ao?

OpenStudy (aaronq):

since they're giving you percentages you can take A to be 12.5 and Ao to be 100

OpenStudy (anonymous):

Okay, I will try that.

OpenStudy (anonymous):

And what is t?

OpenStudy (aaronq):

time elapsed

OpenStudy (aaronq):

what you're solving for in the problem

OpenStudy (anonymous):

Right, I am blushing ;)

OpenStudy (aaronq):

hahah oh you

OpenStudy (anonymous):

I am still unable to get a value that seems appropriate as a possible answer to this

OpenStudy (aaronq):

can you post the steps?

OpenStudy (anonymous):

12.5 = 100 e^0.5545t Have I set it up wrong?

OpenStudy (aaronq):

12.5 = 100 e^(-0.5545t)

OpenStudy (aaronq):

you were only missing the negative sign

OpenStudy (anonymous):

Okay...if I solve the left side, I get 57.435.

OpenStudy (aaronq):

um post each step, because thats not right :S

OpenStudy (anonymous):

Perhaps it's the "e" that is causing me problems...

OpenStudy (aaronq):

do you know how to solve logarithms?

OpenStudy (anonymous):

Natural base e.

OpenStudy (aaronq):

okay, so you have: 12.5 = 100 e^(-0.5545t) what's next?

OpenStudy (anonymous):

I calculated the right side, and i keep getting 57.43

OpenStudy (aaronq):

show me each step

OpenStudy (anonymous):

I type this: 100e^(-0.5545), which gives 57.43. ?

OpenStudy (aaronq):

you can't solve that though because there is a "t" in the exponent

OpenStudy (anonymous):

So, i have to isolate t.

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

Divide 12.5 by 100e?

OpenStudy (aaronq):

just 100, e is to the exponent -0.5545t

OpenStudy (anonymous):

so 12.5/100 = e^(-0.5545)

OpenStudy (anonymous):

t

OpenStudy (aaronq):

okay, whats next?

OpenStudy (anonymous):

multiply e by -0.5545?

OpenStudy (aaronq):

you can't do that because you still have the t

OpenStudy (anonymous):

I have no idea what to do now.

OpenStudy (aaronq):

i thought you said you knew how to solve logarithms :P

OpenStudy (aaronq):

so you have this: 12.5/100 = e^(-0.5545t) "take" ln of both sides, ln(12.5/100) = lne^(-0.5545t) ( you know (or you should know lol) that, lne=1 ) using logarithmic rules take the exponent to the front ln(12.5/100) = lne^(-0.5545t) -> ln(12.5/100) = (-0.5545t)lne since lne =1, so ln(12.5/100) = (-0.5545t)(1) then divide by -0.5545 ln(12.5/100)(-1/0.5545) = t

OpenStudy (aaronq):

do you get it?

OpenStudy (anonymous):

Yes, thank you so much. I appreciate your time and help and patience so much.

OpenStudy (aaronq):

no problem ! i hope you can do them by yourself now (not trying to be rude)